3. (9 Marks) The length of time required by students to complete a coe-hour exam is a random variable with a probability dens i need clear w3. (9 Marks) The length of time required by students to complete a one-bour exam is random variable with a probability densitriting
Answer #1


f(x) = (a tor? rolne for P.d.f. S files.de, at S cotx? dx = 1 [c + 3 اب 1 2 3 - COM = 2.67 O Plosesos) B 8 2 x²+x? da 0 8 20Elx) Sx. fb). dx 20 1을 x 34 12).da 8 !). 1 + 32+ 15 бо 0.73 12 vlx)= Elx?)- (E\x) Elx2= Sx? f(x) dx for? (2 x 32 12).do 0 8